OpenStudy (anonymous):

I need help!! Suppose that S and T are points on the number line. If ST=17 and S lies at -4, where could T be located?

OpenStudy (anonymous):

@cj49

OpenStudy (anonymous):

it wasnt right

OpenStudy (cj49):

t= -4 +17=13 or -4-17=-21